OpenStudy (anonymous):

1/2x+5=1/3x+7 @surdawi

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }x + 5 =\frac{1}{3}x +7\]

OpenStudy (anonymous):

yeap thats it

OpenStudy (anonymous):

lcd is 6; now what u gonna do

OpenStudy (anonymous):

times 6 to each side

OpenStudy (anonymous):

correct do that

OpenStudy (anonymous):

3x+30=2x+36

OpenStudy (anonymous):

oops i mean 42